OpenStudy (anonymous):

A cube has a surface area of one hundred fifty square yards. What is the length of a side? A..five yards B..twelve point five yards C..twenty-five yards D..thirty-seven point five yards

OpenStudy (anonymous):

and if u can,i would LOVE a explanation on how u got it

OpenStudy (anonymous):

150/6: 6 sides in a cube 25 = area of one side side = x^2 = 25 5^2 = 25 one side is 5 A

OpenStudy (anonymous):

thx @rdr006